a mother earned $16250 from royalties on her cookbook. She set aside 20% of this for a down payment for a new home. The balance will be used for her son's future education. She invested a portion of the money in a bank certificate of a deposit (cd account) that earns 4% and the remainder in a savings bond that earns 7%. IF the total interest earned after one year is $780.00, how much money was invested at each rate?
Answer by Boreal(15235)

16,250 *0.20=3250
She has $13,000 left
cd=x
bond=13,000-x
.04x + .07(13,000-x)=780
.04x + 910 -.07x =780
-.03x = -130

x=$4333.33 ;;cd

13000-x=$8666.67 ;;bond
Check:
4333.33*.04=173.32
8666.67*.07=606.67. Sum is $780.